How many ounces of a 35% alcohol solution and a 40% alcohol solution must be combined to obtain 80 ounces of a 37% solution?
Answer by ankor@dixie-net.com(22740)

How many ounces of a 35% alcohol solution and a 40% alcohol solution must be combined to obtain 80 ounces of a 37% solution?
:
let x = amt of 40% solution required
the total is 80 oz, therefore
(80-x) = amt of 35%
:
.40x + .35(80-x) = .37(80)
.40x + 28 - .35x = 29.6
.40x - .35x = 29.6 - 28
.05x = 1.6
x = 1.6/.05
x = 32 oz of the 40% solution
:
I'll let you find amt of 35% solution
Check your solutions in the original equation